# Identifying Causal Direction via Variational Bayesian Compression

### v3

@@ Abstract
Telling apart the cause and effect between two random variables with purely observational data is a challenging problem that finds applications in various scientific disciplines. A key principle utilized in this task is the algorithmic Markov condition, which postulates that the joint distribution, when factorized according to the causal direction, yields a more succinct codelength compared to the anti-causal direction. Previous approaches approximate these codelengths by relying on simple functions or Gaussian processes (GPs) with easily evaluable complexity, compromising between model fitness and computational complexity. To overcome these limitations, we propose leveraging the variational Bayesian learning of neural networks as an interpretation of the codelengths. Consequently, we can enhance the model fitness while promoting the succinctness of the codelengths, while avoiding the significant computational complexity of the GP-based approaches. Extensive experiments on both synthetic and real-world benchmarks in cause-effect identification demonstrate the effectiveness of our proposed method, surpassing the overall performance of related complexity-based and structural causal model regression-based approaches.
@@ 1 Introduction
To address the limitations of these previous compression-based methods, we propose the leverage of neural networks for learning the conditional models, which are considered as universal approximators (Hornik et al., 1989). Moreover, the algorithmic complexity of the networks can be approximated empirically via the concepts of bits-back coding (Hinton & van Camp, 1993; Wallace, 1990) and variational Bayesian coding (Honkela & Valpola, 2004; Louizos et al., 2017; Blier & Ollivier, 2018). In this study, we introduce COMIC—a Bayesian COMpression-based approach to Identifying the Causal direction—that improves the model fitness without compromising the computability of model complexity and avoids the higher computational complexity of GP-based modeling. Correspondingly, through extensive empirical evaluation on benchmarks for the cause-effect identification task, our approach not only outperforms complexity-based methods but also surpasses other related regression-based ones.
@@
    We propose the utilization of Bayesian neural networks for modeling the conditional distributions to address the challenge of balancing between the flexibility and scalability, which is hindering previous complexity/compression-based methods. By minimizing the variational Bayesian codelength, we can approximate the algorithmic complexity of neural networks.
@@
    The capability of our proposed approach is assessed on both synthetic and real-world bivariate causal discovery benchmarks. In comparison with related approaches, our method achieves significant performance improvements in a majority of the benchmarks, demonstrating its effectiveness and consistency in evaluating the empirical complexity of the causal models.
@@ 3 Preliminaries
where $K\left(\cdot\right)$ is the Kolmogorov complexity, $\mathrm{PA}_{G,j}$ denotes the parents of $X_{j}$ in the causal hypothesis $G$, and $\stackrel{{\scriptstyle+}}{{=}}$ denotes equality up to a constant, which is independent of $p\left(\cdot\right)$.
@@ 4.2 Variational Bayesian Code for Evaluating Complexity of Neural Networks
The problem of encoding a conditional distribution $p\left(Y\mid X\right)$ is often defined via a transmitting perspective. Alice has a dataset $\mathcal{D}^{N}\coloneqq\left\{\left(x^{\left(i\right)},y^{\left(i\right)}\right)\right\}_{i=1}^{N}$ that needs to be transported to Bob, who has already got the input samples of this dataset $\left\{x^{\left(i\right)}\right\}_{i=1}^{N}$. The most efficient method is to encode the conditional distribution $p\left(Y\mid X\right)$ so that Bob can predict the remaining output part $\left\{y^{\left(i\right)}\right\}_{i=1}^{N}$ of $\mathcal{D}^{N}$. The variational Bayesian code is a two-part code in MDL where both Alice and Bob first designate a class of model =$\mathcal{M}=\left\{p\left(y\mid x,\boldsymbol{\theta}\right)\mid\boldsymbol{\theta}\in\boldsymbol{\Theta}\right\}$, where $\boldsymbol{\theta}$ represents the parameters of the conditional probability density function $p\left(y\mid x,\boldsymbol{\theta}\right)$, and a prior distribution for the parameters with the probability density function $p\left(\boldsymbol{\theta}\right)$. The corresponding two-part codelength [^2] In most information-theoretic literature, the amount of information is measured in bits. As a result, the logarithm with the base of $2$ is commonly chosen. In this work, for convenience, we will use the natural logarithm and measure the amount of information in nats (natural units of information, Grünwald, 2007). The amount in nats can easily be converted to the corresponding value in bits by dividing it by $\log 2$. for this setting can be computed as
@@
where the first term corresponds to the fitness of the data given the model and the second term corresponds to the complexity of the model of the two-part MDL principle (Honkela & Valpola, 2004; Louizos et al., 2017). $-L_{q_{\phi}\left(\boldsymbol{\theta}\right)}^{\textrm{var}}\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)}\right)$ is known as the evidence lower bound (ELBO) in the variational inference problem (Blei et al., 2017). From this perspective, by minimizing Eq. (10), we can expect to achieve the negative logarithm of the evidence in the Bayesian inference problem, which is also known as the Bayesian codelength in MDL coding (Grünwald, 2007) as follows
@@
[…]
@@ 4.4 Causal Identifiability
Let the two Bayesian causal models $M_{X\rightarrow Y}$ and $M_{Y\rightarrow X}$ respectively be factorized into the marginal densities—$p\left(x\mid M_{X\rightarrow Y}\right)$ and $p\left(y\mid M_{Y\rightarrow X}\right)$—and the conditional densities—$p\left(y\mid x,\boldsymbol{\theta}_{Y},M_{X\rightarrow Y}\right)$ and $p\left(x\mid y,\boldsymbol{\theta}_{X},M_{Y\rightarrow X}\right)$. The marginal densities are standard Gaussian $\mathcal{N}\left(\cdot\mid 0,1\right)$. The conditional densities are Gaussian likelihoods $\mathcal{N}\left(\cdot\mid\mu\left(\cdot;\boldsymbol{\theta}\right),\sigma^{2}\left(\cdot;\boldsymbol{\theta}\right)\right)$, where $\mu\left(\cdot;\boldsymbol{\theta}\right)$ and $\sigma\left(\cdot;\boldsymbol{\theta}\right)$ are parametrized by neural networks as described in Eq. (14), Sec. 4.3, and the parameters $\boldsymbol{\theta}$ follow Gaussian priors specified in App. B. Under these conditions, the Bayesian causal models are not separable-compatible.
@@ 5 Experiments
(solid: HSIC, dashed: likelihood)
(h) CAM & RESIT
(solid: CAM, dashed: RESIT)
Figure 1: Performance on synthetic and real-world benchmarks. Higher accuracy and bidirectional AUROC (Bi-AUROC) values are preferable. Our COMIC method is compared against Sloppy (with the AIC and BIC variants, Marx & Vreeken, 2019a), Sloper (Marx & Vreeken, 2019b) and Slope (Marx & Vreeken, 2017), QCCD (Tagasovska et al., 2020), IGCI (with uniform and Gaussian reference measures, Daniušis et al., 2010), GPLVM (Dhir et al., 2024a), LOCI (with HSIC and maximum likelihood, Immer et al., 2023), CAM (Bühlmann et al., 2014), and RESIT (Peters et al., 2014). COMIC achieves significantly better and more consistent results across all datasets, with a nearly ideal Bi-AUROC on every benchmark. Our promising Bi-AUROC scores suggest strong probabilities that COMIC can correctly determine causal directions on both synthetic and real-world data. In contrast, the baseline methods cannot accomplish satisfactory results when their assumptions are not fulfilled.
Throughout this section, the empirical performance of our COMIC approach [^3] The Python implementation of our method is available at https://github.com/quangdzuytran/COMIC. is evaluated in comparison with state-of-the-art complexity-based and regression-based methods for bivariate causal discovery. Implementation details, descriptions of the benchmarks, and detailed results related to this section are provided in App. DE, and F, respectively. In addition, we conduct a series of ablation studies to assess the contribution of location-scale estimation, the importance of considering model complexity, the effects of varying layer widths and numbers of layers of the neural networks, the impact of employing a sparsity-inducing prior, and alternative modeling choices for encoding the cause, as presented in App. G.
@@ 5.1 Experimental Settings
We compare our work against various baseline ICM-based and FCM-based methods. Methods based on the principle of ICMs include Sloppy (Marx & Vreeken, 2019a), Slope (Marx & Vreeken, 2017, 2019b), QCCD (Tagasovska et al., 2020), and IGCI (Daniušis et al., 2010), and GPLVM (Dhir et al., 2024a). For FCM-based methods, we choose CAM (Bühlmann et al., 2014) and RESIT (Peters et al., 2014) as representatives for ANM-based methods, and LOCI (Immer et al., 2023) to represent LSNM-based methods. Some of these methods have different variants of scoring approaches, which we also include in the evaluations.
@@ 5.2 Results & Discussion
The experimental results for all aforementioned benchmarks are visualized in Fig. 1.
@@
In general, our COMIC approach surpasses the baseline methods on most synthetic benchmarks. Notably, our acquired Bi-AUROC scores are approaching the ideal scores. Higher Bi-AUROC values when compared to corresponding accuracy values can be due to higher confidence scores in correctly predicted cases and very low confidence scores in incorrectly predicted cases.
On the simpler synthetic benchmarks—AN, AN-s, LS, LS-s, and MN-U—COMIC can perfectly determine the causal directions in every case. This favorable outcome may be attributed to the location-scale modeling in Eq. (13), as QCCD, GPLVM and LOCI also achieve comparable scores in accuracy and Bi-AUROC. IGCI with Gaussian reference measure is another method that excels in this group of datasets. However, since the remaining approaches are not designed for data with heteroscedastic noises, they do not perform as well on the more challenging LS, LS-s, and MN-U sets. Particularly, the predictions of Sloper, Slope, IGCI with uniform reference measure, CAM, and RESIT on these datasets are suboptimal compared to other baselines.
Our approach also demonstrates more enhanced and consistent performance on more complicated datasets. On SIM and, especially, SIM-c, FCM-based methods with post-hoc independence testing, such as LOCI with HSIC and RESIT, attain decently high prediction scores. Although COMIC is not explicitly designed for pairs with hidden confounders, our approach demonstrates its potential on SIM-c, which involves a hidden confounder in each pair. Most methods perform decently on SIM-G and SIM-ln, which are more manageable compared to the previous two sets. One notable outlier among the baseline models is IGCI, which exhibits the lowest accuracy and Bi-AUROC on SIM-ln, despite its intended focus on telling apart causal directions in low-noise scenarios. Because the Gaussian noises emerge in assumptions of various methods, including Gaussian-referenced IGCI, LOCI, CAM, and RESIT, they benefit from the configuration of the SIM-G set and exhibit better predictions.
The diverse causal modeling of CE-Multi leads to the divergence in results from different benchmarks. Information-theoretic approaches, including COMIC, Sloppy with BIC, Sloper, Slope, IGCI with Gaussian reference measure, and GPLVM, tend to perform well on this benchmark. QCCD and uniform-referenced IGCI, while also belonging to this category, do not perform as effectively. FCM-based methods designed for additive noise models, such as CAM and RESIT, also exhibit inefficacious performance on this dataset. A majority of methods featuring regressions in the learning process can adequately predict causal relations of the CE-Net benchmark. IGCI is the only baseline tested that underperforms on this benchmark. Due to the difficulty of CE-Cha, the results on this set are quite similar to SIM, where most baselines struggle, except for COMIC, GPLVM, LOCI with HSIC, and RESIT. This discrepancy highlights the advantage of the greater modeling flexibility offered by advanced regression techniques in COMIC and GPLVM in comparison to other complexity-based approaches.
@@
On the Tübingen benchmark, it is obvious that the identification ability of COMIC is considerably preferable compared to the chosen baselines. Similar to the synthetic benchmarks, the Bi-AUROC in this real-world dataset is also approximating the ideal value, indicating the high probability of correctly predicting the causal relation when the confidence is high. As noted by Marx & Vreeken (2019a), Sloppy with AIC offers greater flexibility for more complex datasets, such as this dataset, and performs better than the BIC variant. This behavior also appears in IGCI where the variant with the uniform reference measure obtains higher accuracy and Bi-AUROC on this dataset. Although Sloper and LOCI with HSIC are better than their respective variants on previous benchmarks, Slope and LOCI with maximum likelihood, they are outperformed by these variants on this benchmark. GPLVM, leveraging its latent variable modeling, acquires a comparable Bi-AUROC score to other compression-based methods including Sloppy with AIC, Slope, and QCCD on this dataset. This also showcases the effectiveness of these compression-based methods, whose Bi-AUROC scores are notably higher than the remaining baseline methods. Besides our COMIC approach, QCCD is the method with the second-highest accuracy score.
@@ 6 Conclusion
In this work, we have proposed COMIC—a neural network compression-based approach for determining the cause and effect via the variational Bayesian code—where a more universal and scalable class of neural networks is utilized for modeling the conditionals to improve fitness and induce better codelengths. With the variational Bayesian coding scheme, the algorithmic complexity of these networks can be assessed empirically to approximate the theoretical Kolmogorov complexity, whose identifiability can also be scrutinized in relation to the marginal likelihood-based perspective. The effectiveness of COMIC has been validated through comprehensive experiments across various benchmark datasets, delivering promising results and demonstrating superior performance compared to related methods based on the compression and FCM regression objectives.
@@
Acknowledgements
This research was partially supported by the Wellcome Trust [grant number: 303030/Z/23/Z].
@@ Appendix A Kolmogorov Complexity, Algorithmic Mutual Information, & Minimum Description Length
where $x^{*}$ is the shortest description of the string $x$ and $\stackrel{{\scriptstyle+}}{{=}}$ denotes equality up to an additive constant, which only depends on $\mathcal{U}$ and is independent of $x$ and $y$.

### v4

@@ Abstract
Telling apart the cause and effect between two random variables with purely observational data is a challenging problem that finds applications in various scientific disciplines. A key principle utilized in this task is the algorithmic Markov condition, which postulates that the joint distribution, when factorized according to the causal direction, yields a more succinct codelength compared to the anti-causal direction. Previous approaches approximate these codelengths by relying on simple functions or Gaussian processes (GPs) with easily evaluable complexity, compromising between model fitness and computational complexity. To address these limitations, we propose leveraging the variational Bayesian learning of neural networks as an interpretation of the codelengths. This allows the improvement of model fitness, while maintaining the succinctness of the codelengths, and the avoidance of the significant computational complexity of the GP-based approaches. Extensive experiments on both synthetic and real-world benchmarks in cause-effect identification demonstrate the effectiveness of our proposed method, showing promising performance enhancements on several datasets in comparison to most related methods.
@@ 1 Introduction
To address the limitations of these previous compression-based methods, we propose the leverage of neural networks for learning the conditional models, which are regarded as universal approximators (Hornik et al., 1989). Moreover, the algorithmic complexity of the networks can be approximated empirically via the concepts of bits-back coding (Hinton & van Camp, 1993; Wallace, 1990) and variational Bayesian coding (Honkela & Valpola, 2004; Louizos et al., 2017; Blier & Ollivier, 2018). In this study, we introduce COMIC—a Bayesian COMpression-based approach to Identifying the Causal direction—that improves the model fitness of compression-based methods without compromising the computability of model complexity and avoids the higher computational complexity of GP-based modeling. Correspondingly, through extensive empirical evaluation on benchmarks for the cause-effect identification task, our approach demonstrates promising results with performance improvements compared to a majority of complexity-based and maximum likelihood-based approaches on multiple benchmarks.
@@
    We propose the utilization of Bayesian neural networks for modeling the conditional distributions to address the challenge of balancing flexibility and scalability, which is hindering previous complexity/compression-based methods. By minimizing the variational Bayesian codelength, we can approximate the algorithmic complexity of neural networks.
@@
    The capability of our approach is assessed on both synthetic and real-world bivariate causal discovery benchmarks. In comparison to most related approaches, our method achieves performance improvements in several benchmarks, demonstrating the effectiveness of our approach in identifying the causal direction.
@@ 3 Preliminaries
[…]
@@ 4.2 Variational Bayesian Code for Evaluating Complexity of Neural Networks
The problem of encoding a conditional distribution $p\left(Y\mid X\right)$ is often defined via a transmitting perspective. Alice has a dataset $\mathcal{D}^{N}\coloneqq\left\{\left(x^{\left(i\right)},y^{\left(i\right)}\right)\right\}_{i=1}^{N}$ that needs to be transported to Bob, who has already got the input samples of this dataset $\left\{x^{\left(i\right)}\right\}_{i=1}^{N}$. The most efficient method is to encode the conditional distribution $p\left(Y\mid X\right)$ so that Bob can predict the remaining output part $\left\{y^{\left(i\right)}\right\}_{i=1}^{N}$ of $\mathcal{D}^{N}$. The variational Bayesian code is a two-part code in MDL where both Alice and Bob first designate a class of model $\mathcal{M}=\left\{p\left(y\mid x,\boldsymbol{\theta}\right)\mid\boldsymbol{\theta}\in\boldsymbol{\Theta}\right\}$, where $\boldsymbol{\theta}$ represents the parameters of the conditional probability density function $p\left(y\mid x,\boldsymbol{\theta}\right)$, and a prior distribution for the parameters with the probability density function $p\left(\boldsymbol{\theta}\right)$. The corresponding two-part codelength [^2] In most information-theoretic literature, the amount of information is measured in bits. As a result, the logarithm with the base of $2$ is commonly chosen. In this work, for convenience, we will use the natural logarithm and measure the amount of information in nats (natural units of information, Grünwald, 2007). The amount in nats can easily be converted to the corresponding value in bits by dividing it by $\log 2$. for this setting can be computed as
@@
where the first term corresponds to the fitness of the data given the model and the second term corresponds to the complexity of the model of the two-part MDL principle (Honkela & Valpola, 2004; Louizos et al., 2017). $L_{q_{\phi}\left(\boldsymbol{\theta}\right)}^{\textrm{var}}\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)}\right)$ is known as the evidence lower bound (ELBO) in the variational inference problem (Blei et al., 2017). From this perspective, by minimizing Eq. (10), we can expect to achieve the negative logarithm of the evidence in the Bayesian inference problem, which is also known as the Bayesian codelength in MDL coding (Grünwald, 2007) as follows
@@
$\coloneqq-\log\int p\left(y^{\left(1:N\right)}\mid x^{\left(1:N\right)},\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)d\boldsymbol{\theta}.$ (11)
@@ 4.4 Causal Identifiability
[…]
@@ 5 Experiments
(h) CAM
Figure 1: Performance on synthetic and real-world benchmarks. Higher accuracy and bidirectional AUROC (Bi-AUROC) values are preferable. Our COMIC method is compared against Sloppy (with the AIC and BIC variants, Marx & Vreeken, 2019a), Sloper (Marx & Vreeken, 2019b) and Slope (Marx & Vreeken, 2017), QCCD (Tagasovska et al., 2020), IGCI (with uniform and Gaussian reference measures, Daniušis et al., 2010), GPLVM (Dhir et al., 2024a), LOCI (Immer et al., 2023), and CAM (Bühlmann et al., 2014). The marginal likelihood-based objective in COMIC achieves promising results with enhanced performance compared to methods with the maximum likelihood-based objectives in LOCI and CAM, especially on the real-world Tübingen benchmark. In comparison to most complexity-based methods, COMIC also achieves better performance on multiple synthetic datasets and obtains comparable results on real-world data.
Figure 2: Overall performance of COMIC and the baseline methods on all benchmarks. The accuracy and bidirectional AUROC (Bi-AUROC) scores are averaged over all datasets. The results indicate that our COMIC approach achieves better and more consistent overall results in comparison to a majority of baselines.
Throughout this section, the empirical performance of our COMIC approach is evaluated in comparison with state-of-the-art complexity-based and regression-based methods for bivariate causal discovery. Implementation details and descriptions of the benchmarks related to this section are provided in App. D and E, respectively.
@@ 5.1 Experimental Settings
We compare our work against various baseline ICM-based and FCM-based methods. Methods based on the principle of ICMs include Sloppy (Marx & Vreeken, 2019a), Slope (Marx & Vreeken, 2017, 2019b), QCCD (Tagasovska et al., 2020), and IGCI (Daniušis et al., 2010), and GPLVM (Dhir et al., 2024a). For FCM-based methods, we choose CAM (Bühlmann et al., 2014) as a representative for ANM-based methods, and LOCI (Immer et al., 2023) with maximum likelihood scoring to represent LSNM-based methods. Some of the ICM-based methods have different variants, which we also include in the evaluations.
@@ 5.2 Results & Discussion
The experimental results on all aforementioned benchmarks are visualized in Fig. 1 and 2. The overall results in Fig. 2 demonstrate that our COMIC approach achieves second-best performance in comparison to complexity-based and maximum likelihood-based methods. Our method is also more consistent across all benchmarks, as indicated by the smaller standard deviation of the accuracy and Bi-AUROC scores compared to most baseline methods. Furthermore, COMIC outperforms the maximum likelihood-based methods, i.e., LOCI and CAM, suggesting that its marginal likelihood-based objective delivers promising results with improved performance relative to methods with the maximum likelihood-based objectives.
GPLVM is the baseline that attains the best overall results. However, it is important to note that GPLVM also requires the most substantial computational resources, with execution on GPUs (as noted in App. D), compared to COMIC and other baselines, which can be executed on CPUs. This high resource demand stems from the computational complexity of Gaussian processes with latent variables and the extensive use of random restarts for hyperparameter selection.
@@
On synthetic AN, AN-s, LS, LS-s, and MN-U benchmarks, COMIC perfectly determines the causal directions in every case. QCCD, GPLVM, and LOCI also achieve equivalent accuracy and Bi-AUROC scores on these datasets. These findings align with theoretical expectation that models identifiable via maximum likelihood are also identifiable via marginal likelihood (Dhir et al., 2024a). IGCI with Gaussian reference measure is another method that excels in this group of datasets. The remaining baseline approaches do not perform as well on the more challenging LS, LS-s, and MN-U sets. In particular, the predictions of Sloper, Slope, IGCI with uniform reference measure, and CAM on these datasets are noticeably suboptimal compared to other baselines.
On the more challenging SIM and SIM-c datasets, our COMIC approach demonstrates performance comparable to complexity-based methods, such as Sloppy with AIC, Sloper, and Slope. A majority of baseline methods perform decently on SIM-G and SIM-ln, which are more manageable compared to the previous two sets. COMIC achieves the second-highest scores on SIM-ln, with slightly lower accuracy and Bi-AUROC than GPLVM. One unexpected outlier among the baseline models is IGCI, which exhibits the lowest accuracy and Bi-AUROC on SIM-ln, despite its intended focus on telling apart causal directions in low-noise scenarios. Moreover, Sloppy with AIC, Sloper, Slope, and QCCD underperform relative to COMIC on the SIM-G dataset with nearly Gaussian distributions over the causes.
The diverse causal modeling of CE-Multi results in varying performance across benchmarks. Information-theoretic approaches, including COMIC, Sloppy with BIC, Sloper, Slope, IGCI with Gaussian reference measure, and GPLVM, tend to perform well on this benchmark. QCCD and uniform-referenced IGCI, while also belonging to this category, perform less effectively. Since CAM is designed for additive noise models, it exhibits inefficacious performance on this dataset. A majority of methods featuring regressions in the learning process can adequately predict causal relations of the CE-Net benchmark. Hence, IGCI is the only baseline tested that underperforms on this benchmark. Due to the difficulty of CE-Cha, the results on this set resemble those on SIM, where most methods struggle, except for GPLVM.
@@
On the real-world Tübingen benchmark, complexity-based methods, including our COMIC approach, GPLVM, Sloppy with AIC, Slope, and QCCD, achieve comparable accuracy and Bi-AUROC scores. This highlights the effectiveness of compression-based methods, whose Bi-AUROC scores are notably higher than the remaining maximum likelihood-based methods that focus solely on model fitness. As noted by Marx & Vreeken (2019a), Sloppy with AIC offers greater flexibility for more complex datasets, such as this one, and performs better than the BIC variant. A similar pattern appears in IGCI, where the uniform reference measure variant obtains higher accuracy and Bi-AUROC. Although Sloper outperforms its respective variant Slope on previous benchmarks, it is surpassed by Slope on this benchmark.
@@ 6 Conclusion
In this work, we have proposed COMIC—a neural network compression-based approach for determining the cause and effect via the variational Bayesian code—where a more universal and scalable class of neural networks is utilized for modeling the conditionals to improve fitness and induce better codelengths. With the variational Bayesian coding scheme, the algorithmic complexity of these networks can be assessed empirically to approximate the theoretical Kolmogorov complexity, and its identifiability can also be scrutinized from a marginal likelihood-based perspective. The effectiveness of COMIC has been validated through comprehensive experiments, delivering promising results and demonstrating enhanced performance compared to most related methods based on the compression and FCM regression objectives across multiple benchmarks.
@@ Appendix A Kolmogorov Complexity, Algorithmic Mutual Information, & Minimum Description Length
[…]